OpenStudy (anonymous):

the width of a rectangle is 2 centimeters less than one-half its length. if the perimeter is 56 centimeters, find the length and width.

OpenStudy (anonymous):

2w+2l=56w=0.5l-2

OpenStudy (anonymous):

ignore that

OpenStudy (anonymous):

with the equations 2w+2l=56 w-0.5l=-2 create a 2x2 matrix A=\[\left[\begin{matrix}2 & 2 \\ 1 & -0.5\end{matrix}\right]\] with an answer matrix of \[\left(\begin{matrix}56 \\ -2\end{matrix}\right)\] create two more matrices switching a row of the first matrix with the answer matrix B=\[\left[\begin{matrix}56 & 2 \\ -2 & -0.5\end{matrix}\right]\] C=\[\left[\begin{matrix}2 & 56 \\ 1 & -2\end{matrix}\right]\] the w=det(B)/det(A) and l=det(C)/det(A) w=5(6(-0.5)-(-2)(2))/((2)(-0.5)-2)= -24/-3 w=8 l=((2)(-2)-56))/-3=-60/-3 l=20

OpenStudy (anonymous):

I hope I helped

OpenStudy (anonymous):

I was actually laughing the whole time I wrote that